# Soft Prompting for Unlearning in Large Language Models

### v1

@@ 5 Conclusion
In this work, we investigate unlearning in LLMs to remove the influence of unwanted training examples during text classification. We present a soft prompting strategy to unlearn subsets of training data while keeping pre-trained LLM parameters frozen to maintain the model’s generalizability. Our SPUL framework optimizes a small number of prompt tokens using a multi-objective loss function defined on disjoint training data subsets representing the forget data that is subjected to removal and the retain data that aims to preserve model utility. Experimental evaluation on sentiment classification datasets demonstrates the superior efficiency of our soft prompting-based unlearning over fine-tuning-based baselines. We also empirically show that SPUL can adapt to multiple LLMs and is robust to a high number of unlearning samples.
@@ Acknowledgements
We address the limitations of this work in the following. Our experiments primarily focus on open-source LLMs as the soft prompting framework requires access to frozen pre-trained parameters to compute gradients for the soft prompts despite not needing to update the LLM parameters. Furthermore, this work focuses on the task of text classification, specifically sentiment classification for the formulation of the unlearning framework and evaluation. Future research could explore the efficiency of soft prompting to achieve unlearning in the context of NLP tasks such as text generation, question answering, text summarization, and so on. Also, the soft prompting unlearning framework has not been evaluated comprehensively as we emphasize performance metrics to demonstrate unlearning efficacy. We note that there is a lack of an extensive evaluation pipeline for LLM unlearning in the current literature. Further research is needed to evaluate the robustness of the framework subject to model-stealing attacks, MIAs, and jailbreaking attempts.
@@
In this study, our focus is to achieve LLM unlearning in a resource-efficient manner. We aim to enable forgetting of unwanted and undesirable knowledge as per users’ requests while maintaining model efficiency to avoid exploitation of protected information. The datasets used for evaluation are publicly available and implemented within the intended use. Our usage of publicly available pre-trained LLMs also adheres to the associated licenses. We hope our study can further the research and literature on resource-efficient LLM unlearning.

### v2

@@ 5 Conclusion
In this work, we investigate unlearning in LLMs to remove the influence of unwanted training examples during text classification and multiple-choice question answering. We present a soft prompting strategy to unlearn subsets of training data while keeping pre-trained LLM parameters frozen to maintain the model’s generalizability. Our SPUL framework optimizes a small number of prompt tokens using a multi-objective loss function defined on disjoint training data subsets representing the forget data that is subjected to removal and the retain data that aims to preserve model utility. Experimental evaluation on sentiment classification and QA datasets demonstrates the superior efficiency of our soft prompting-based unlearning over fine-tuning-based baselines. We also empirically show that SPUL can adapt to multiple LLMs and is robust to a high number of unlearning samples.
@@ Acknowledgements
We address the limitations of this work in the following. Our experiments primarily focus on open-source LLMs as the soft prompting framework requires access to frozen pre-trained parameters to compute gradients for the soft prompts despite not needing to update the LLM parameters. Furthermore, this work focuses on the task of text classification, specifically sentiment classification, and question answering for the formulation of the unlearning framework and evaluation. Future research could explore the efficiency of soft prompting to achieve unlearning in the context of NLP tasks such as text generation, text summarization, and so on. Also, the soft prompting unlearning framework has not been evaluated comprehensively as we emphasize performance metrics to demonstrate unlearning efficacy. We note that there is a lack of an extensive evaluation pipeline for LLM unlearning in the current literature. Further research is needed to evaluate the robustness of the framework subject to model-stealing attacks, MIAs, and jailbreaking attempts.
@@
In this study, our focus is to achieve LLM unlearning in a resource-efficient manner. We aim to enable forgetting of unwanted and undesirable knowledge as per users’ requests while maintaining model efficiency to avoid exploitation of sensitive information. The datasets used for evaluation are publicly available and implemented within the intended use. Our usage of publicly available pre-trained LLMs also adheres to the associated licenses. We hope our study can further the research and literature on resource-efficient LLM unlearning.